Integrated Nanosystems Research Facility

 

Equipment:

  • Facilities is going to perform a sanitation and cleaning procedure in the DI water system Wednesday, September 5th @ 6 am. This procedure will take three days and water will not be available until September 10th.
  • The E-Beam evaporator is still being worked on and we have sent out various components for bead blasting. The items were picked up Tuesday and they need three working days to clean. The system should be back online by September 10th.
  • Oriel UV Flood exposure is being looked at. We have ordered some replacement components in the high voltage portion of the system. We are waiting to receive those parts.
  • K&S 780 dicing saw is back online.

Note:
Wet bench area – Due to the exessive amout of users using the wet etch area, users have requested reserving time to use the benches. The idea is to mount a white board in the robing area for users to reserve certain benches at specific times. Hopefully this will reduce the congestion and make a safer environment. If you have any questions please ask staff.

Thanks,
INRF Staff

Hello Users,

For all new foundry requests, no matter how big or small, please fill out our new online request form.

http://www.inrf.uci.edu/facility/services/service-request/

Thanks,
INRF Staff

Equipment Updates:

  • The E-beam 1 experienced a high voltage problem on start up. The high voltage circuit needs to be repaired due to excessive amount of material deposited inside the chamber. We will clean the system, repair the damaged circuitry, disassemble and correct the high vacuum valve so it will open and close without hesitation. We are hoping the system will be available sometime next week.
  • The Nanoimprinter back on-line. We drained the oil out of the foreline.
  • We are expecting the CCD camera for the K&S 780 dicing saw shortly and the system should be back up by Wednesday.
  • Oreol Flood exposure is still down and we did not have a chance to look at the power supply.
  • The Karl Suess MA-6 is experiencing problem on channel two. The system is presently set at 5 milliwatts. If using system at 10 milliwatts you will need to double your exposure time. Jake will try to look at this system this weekend.
  • The Critical point dryer received the O rings and we are planning a test run before the end of the week. Check website to verify system is back online.

Thanks,
INRF Staff

We were informed that exhaust fan #4 on the roof is down and HVAC is working to get it back up. If you notice anything strange in the cleanroom (hood alarms, ect…) please let staff know right away, we are not sure what this fan is connected to.

Equipment Status:
  • We are still experiencing problems with the K&S 780 dicing saw. We will be sending the camera system and mixing box out for evaluation and repair. We should have it back sometime next week.
  • Still working on the critical point dryer and talking with their service department to determine why its consistently leaking CO2.
  • The Oreole flood exposure is still experiencing a power supply problem.
  • We’ve completed the ALD system and are planning a test run before the end of the week.
  • Tested out wet oxidation system. A short in gas panel has been repaired and the system is working again.
  • Worked on IM milling, replaced the emissions element and are planning a test run to verify system is operating properly.

Safety Topic:
We experienced an incident Tuesday afternoon. A user may have been exposed to HF due to another user improperly blowing dry an object outside the safety hood.
Users are reminded: Any work within the acid rooms should be performed inside the hood and directed away from yourself and others. If a user is not sure of a procedure, contact staff before proceeding. Users be aware, if involved in any safety incident to notify staff immediately no matter how small so we can take action.

Thanks,
INRF Staff

Finished repairing the CDO this weekend. Due to that repair the STS, Plasma therm and BMR low temp PECVD are back online.

Finished repairs on Hitachi SEM. This system is back online.

We worked on the IM milling, identified a vacuum leak and also replaced the neturalization filament. We are still experiencing problems with the emmision current. We will be addressing that.

Experiencing a problem with the critical point dryer. It is leaking at high pressure. We plan to replace the two retaining rings on the chamber.

We were informed Wednesday that the flood exposure in the INRF UV lamp was not functioning. We recommend if you have a need at this time you utilize the flood exposure in the BiON.

The Thermco 4″ Anneal furnace was contaminated with debris. Furnace tube has been cleaned and is back online.

Information:
Users please use the blog to find out status of lab equipment. As equipment comes back online there are two areas to review for status, our website or the blog. We will updating the blog as equipment is back online real time.

Thank you for your patiences.
INRF Staff

Below is a link to our new online foundry form that is emailed directly to staff. Once submitted you will be contacted by via phone/email.

 

www.inrf.uci.edu/facility/services/service-request/

STS plasma therm RIE & PECVD & low temperature BMR oxide and Nitrite systems are all off line due to our CDO (polution control device) that is required to be in operation while operating these tools experienced a high temperature alarm on the exhaust. While looking into this problem we noticed the reactive gases are eroding internal components. These components must be replaced. There is a good chance the systems will be back online this weekend.

Ion Milling system experienced a chamber vacuum leak and also the neutraliser wire needs to be replaced. At the present time we are using a leak detector to locate and verify the leak.
System should be back online by Monday.

ALD has been disassembled and the chamber and vacuum components to be sent out to be bead blasted. System should be back by Tuesday and we will be reassembled on Wednesday.

Gasonics experienced a faulty replacement part. We received the new part and has been installed but needs to be electronically set up. We hope to have the system up by this weekend.

Hitachi SEM baked out system for last three days. We are replacing the apature and realigning the beam. If all goes well the system should be back up by Thursday evening or Friday.

A Nano imprinter user was using the system Wednesday and noticed oil in the vacuum foreline. The system is tagged out temporily while I clean the foreline. It should be back online by Friday.

Saftey and Observation note.
USERS NEED TO BE MORE CONSCIENTIOUS ABOUT CLOSING WASTE CONTAINERS AFTER COMPLETION OF USING THE CONTAINERS. VERIFY THAT THE STORAGE CONTAINERS ARE FOR REACTIVE MATERIALS, SUCH AS SOLVENT AND ACIDS, AND THEY ARE THE PROPER CONTAINERS AND PROPERLY IDENTIFIED.

We observed while inspecting the lab in the evenings that users are leaving the microscope light on at high intensity. This creates more work for staff replacing lamps. If users do not know how to operate the microscopes properly please contact staff. We also noticed the stage locks being over tightened. Users are forcing the stage up and down which damages the microscope.
Your help in this area would be appreciated.

Thanks,
INRF Staff